A proposed car air bag uses a canister of CO2 gas to inflate a 54 L bag at 20 ∘C and 1.04 atm pressure. What mass of CO2 must the canister contain? Express your answer to two significant figures.
Answer #1

Given:

P = 1.04 atm

V = 54.0 L

T = 20.0 oC

= (20.0+273) K

= 293 K

find number of moles using:

P * V = n*R*T

1.04 atm * 54 L = n * 0.08206 atm.L/mol.K * 293 K

n = 2.336 mol

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

use:

mass of CO2,

m = number of mol * molar mass

= 2.336 mol * 44.01 g/mol

= 1.028*10^2 g

Answer: 1.0*10^2 g